OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 5n + 6 = -4

OpenStudy (anonymous):

5n + 6 = -4 5n = -10 n= -2

OpenStudy (anonymous):

The check: 5(-2) +6=-4 -10+6=-4 -4=-4

OpenStudy (anonymous):

5n + 6 = -4 5n = -10 n=-10/5 n=-2